# SafeNav: Safe Path Navigation using Landmark Based Localization in a GPS-denied Environment

### v1

@@ IV-B3 Position Update
In this model, the next velocity $v_{t+1}$ is computed from the current velocity $v_{t}$, constrained by acceleration, deceleration, and speed limits. Heading $\theta_{t+1}$ is adjusted by $\Delta\theta$ within maneuverability bounds, and the new position $\vec{p}_{t+1}$ is calculated using the updated velocity and heading, assuming linear motion with instantaneous changes.
However, real battlefield dynamics are influenced by unpredictable factors like terrain, obstacles, and environmental changes. To reflect this, the motion model incorporates process noise in position, velocity, acceleration, and heading represented by terrain effects $(\tau)$ and noise terms ${\epsilon}_{x,t},{\epsilon}_{y,t},{\epsilon}_{v_{x},t},{\epsilon}_{v_{y},t},{\epsilon}_{a_{x},t},{\epsilon}_{a_{y},t},{\epsilon}_{\theta,t}$. The updated model is defined in Eqns. 6 and 7.
@@
Terrain effect $\tau$ significantly impacts position prediction accuracy. Accurately estimating $\tau$ improves navigation in complex terrains. The total travel time from point P1 $(x1,y1)$ to P2 $(x2,y2)$, accounting for $\tau$, is given by Eqn. 8.
@@
where $\lambda=\tau/v_{c}$ represents the time delay caused by terrain effect $\tau$, initially estimated using prior knowledge. As terrain and velocity vary, $\tau$ may change during navigation. The predicted position $p_{t+1}$ accounting for $\tau$ is given in Eqn. 9.
@@
where ${p}_{t}$ is the current position of the ME.
@@ IV-C Trajectory Estimation using BMM and EKF
The Extended Kalman Filter (EKF) for the non-linear battlefield scenario linearizes the current mean and covariance, allowing it to approximate the state of a nonlinear system, and thus is an effective filtering algorithm due to its low computational complexity and high estimation accuracy. We adapted an EKF [13] by integrating the BMM defined in Eqns. 6 and 7 in its prediction step to predict the future state of a moving entity.
The state transition function $X_{k+1}$ and the measurement function $Z_{k+1}$ are represented in Eqn. 10. The predicted state estimate $\hat{x}_{k+1}$ and predicted covariance estimate $P_{k+1}$ are given by Eqn. 11. The measurement residual $\tilde{y}_{k+1}$ and residual covariance $S_{k+1}$ are calculated using Eqn. 12.
@@
where, $f(\cdot)$ is the nonlinear state transition function, $U_{k}$ is the control input, and $X_{k}$ is the previous state. $h(\cdot)$ represents the LanBLoc measurement function. $W_{k}$ and $V_{k+1}$ are process and measurement noise, assumed to be normally distributed with covariances $Q_{k}$ and $R_{k+1}$, respectively. $F_{k}$ is the Jacobian of $f(\cdot)$ with respect to $X$, evaluated at $\hat{x}_{k+1}$. $z_{k+1}$ is the observed position at time $k+1$.
The Kalman Gain is a fundamental component of the Kalman Filter that determines the degree to which the new measurement will be incorporated into the state estimate.
So, the Kalman Gain $K_{k+1}$ is computed using the predicted covariance estimate $P_{k+1}$, the transpose of the observation model $H_{k+1}^{T}$, and the inverse of the residual covariance $S_{k+1}^{-1}$ in equation 13 below.
@@ IV-D1 Convex Hull Method
The convex hull method offers a geometrically robust and computationally efficient way to define safe navigation boundaries. In our approach, it verifies whether the predicted position lies within a safe zone while steering toward a look-ahead point. The path is divided into sequential segments, each represented as $S={(x_{1},y_{1}),...,(x_{n},y_{n})}$, and their convex hulls are computed using Chan’s algorithm [10]. The predicted state is then checked using a point-in-polygon test [11].
At each trial’s start, the node’s initial position $(x_{i},y_{i})$ is obtained via LanBLoc, while velocity $(v_{i})$ and heading $(\theta_{i})$ are measured using an IMU sensor simulator. These inputs feed into the EKF module, which predicts the next state $(x_{P},y_{P},v_{P},\theta_{P})$ as detailed in Section IV-B. The node then checks if the predicted position lies within the convex hull of the current or next segment. If so, it proceeds with default control commands; otherwise, it recalculates them or steps back if outside adjacent hulls. Trajectory prediction and safety checks are detailed in Sections IV-C and Algorithm 1, respectively.
@@ IV-D2 Centroid Method
The Centroid Method relies on the centroids of sub-segments within a safe path to compute the control commands. In this approach, the current path segment is divided into smaller sub-segments, and the centroids of these sub-segments are calculated. The algorithm determines the closest centroid to the entity’s current position(updated by EKF) and computes the control input $U(v_{i},{\theta_{i}})$ required to reach the centroid. Using these control commands, it predicts the next position of the entity, which is mentioned in detail in the Algorithm 2. This algorithm attracts the moving entity toward the centroid of the path segment to minimize the drift from the central trajectory.
@@ IV-E1 Rapidly Exploring Random Tree-Star (RRT*)
RRT* [12] finds the shortest obstacle-free path using:
@@
IV-E2 Risk-Aware RRT* (RAw-RRT*)
We extend RRT* to include hazard awareness via a modified cost:
@@ IV-E2 Risk-Aware RRT* (RAw-RRT*)
Here, $C_{\text{length}}$ is the total Euclidean path length, and $C_{\text{risk}}$ sums the Weighted Risk Scores (WRS) across trajectory points. Parameters $\alpha$ and $\beta$ balance between path length and risk exposure. This version is used to re-generate safe paths during CHull and Centroid-based motion.
@@ IV-E3 Edge-Wise WRS Calculation
Deviation from the central path is calculated using Eqn. 17 and WRS is computed based on zone-specific deviation using Eqn. 18.
@@
where $w_{j}$ is the zone weight and $w_{\text{max}}$ applies for large deviations exceeding $d_{\text{max}}$.
IV-E4 Total Trajectory Cost
For a trajectory with points $\{\mathbf{p}_{i},\ldots,\mathbf{p_{n}}\}$, the total risk-aware cost is given by Eqn. 19.
$C_{\text{total}}=\alpha\sum_{i=1}^{n-1}\|\mathbf{p}_{i}-\mathbf{p}_{i+1}\|+\beta\sum_{i=1}^{n}\text{WRS}(\mathbf{p}_{i})$ (19)
@@ IV-F Safe Path Navigation using SafeNav-Chull and SafeNav-Centroid and RAw-RRT*
9:         Use equation 19 to find alternative path to nextPosition that minimize cost.
@@ V Experiments and Results
To test the validity of our proposed navigation algorithms, experiments were performed on an Alienware Aurora R12 System with an 11th Gen Intel Core i7 CPU, 32 GiB of Memory, and an Nvidia GeForce RTX 3070 GPU using Python 3.11 on PyCharm CE 2024 IDE. The landmark recognition model was trained using Python 3.10.12 and torch 2.5.1+cu121, utilizing an NVIDIA A100-SXM4 GPU with 40GB of memory.
@@ V-A Datasets
V-A1 Landmark and Stereo Vision Dataset
We used the publicly available MSTlandmarkv1 dataset to train the landmark recognition model and MSTlandmarkStereov1 for distance estimation from [2].
V-A2 Trajectory Dataset and Safe Path
@@
ii) Real-Imitated Trajectory Dataset: We generated real-imitated trajectories and safe path datasets to validate our navigation framework. A virtual battlefield region of a $200\times 200$ grid was created, mapping landmarks from the MSTlandmarkStereov1 datasets. We randomly placed 25 obstacles and 25 hazard components within this grid to simulate battlefield conditions. Clusters containing at least three landmarks within detectable range were selected as a trilateration set for LanBloc localization. Path clusters, comprising sequences of these landmark clusters, were used to generate ground truth trajectories by navigating through cluster centroids.
iii) Safe Path Buffer and Risk Zone: A safe buffer is a bounded region around the ground truth trajectory, divided into four equal-width zones with risk weights $w_{1}=2$, $w_{2}=4$, $w_{3}=6$, and $w_{4}=8$. The Weighted Risk Score (WRS), calculated using Eqn. 18, increases with distance from the central trajectory, quantifying navigation risk based on deviation.
@@ V-B Evaluation Metrics
We used standard metrics such as ADE, FDE, and Percent Error to evaluate trajectory accuracy, along with a novel AWRS to assess risk exposure. Each safe path was simulated over $1000$ trials and averaged.
The AWRS is defined as the mean of the individual Weighted Risk Scores across all trajectory points: $\text{AWRS}=\frac{1}{N}\sum_{i=1}^{N}\text{WRS}_{i}$, Where N is the total number of trajectory points and $\text{WRS}_{i}$ is the weighted risk score for the $i^{th}$ point calculated using Eqn. 18. ADE is the mean L2 distance between the estimated and ground truth trajectory: $\text{ADE}=\frac{1}{N}\sum_{t=t_{0}}^{t_{f}}\|\hat{Y}_{t}-Y_{t}\|$. FDE is the L2 distance between the estimated and ground truth trajectory at the final time step: $\text{FDE}=\|\hat{Y}_{t_{f}}-Y_{t_{f}}\|$. Percent Error is given by |Est. LengthTrue Length|True $\frac{|\text{Est. Length}-\text{True Length}|}{\text{True Length}}\times 100\%$.
@@ V-C Simulation and Experimental Results Discussion
Baselines:
We evaluated the proposed LanBLoc-BMM against state-of-the-art methods—DeepLoc-GPS, DeepLoc-EKF-GPS [6], and Multi-Sensor Fusion [15]—using both synthetic and real-imitated datasets. Performance with EKF and PF on synthetic and real datasets is summarized in Tables I and II, respectively.
Table I highlights the best overall performance of LanBLoc(EKF), with the lowest ADE, FDE, AWRS, and runtime, while DeepLoc-GPS-EKF $(EKF)$ is a close contender, performing well in all metrics, particularly in FDE and AWRS. DeepLoc-GPS $(EKF)$ also performs well, especially in ADE and FDE; however, it lags in runtime. Sensor Fusion (EKF), while having the highest ADE, shows competitive FDE but is less efficient in computation.
TABLE I: Performance across Algorithms on synthetic trajectory data
@@
From Table II, it is evident that LanBLoc-BMM (EKF) is best regarding accuracy, weighted risk score, and runtime for real trajectory datasets. It consistently offers the most accurate results with reasonable computation times and minimal risk exposure, making it the preferred choice for accurate and efficient localization across different path types. DeepLocGPS(EKF) provides its best accuracy with an ADE/FDE of 0.3041/0.3103, AWRS of 0.2009, and a reasonable runtime of 20.58 ms. DeepLocGPSEKF shows the best accuracy with an ADE/FDE of 0.2869/0.2969, AWRS of 0.1906, and a runtime of 18.94 ms. Sensor Fusion(EKF) provides a balance of accuracy (ADE/FDE of 0.2998/0.3115), AWRS of 0.1749, and an acceptable runtime of 22.31 ms. The proposed LanBLoc-BMM (EKF) based navigation again stands out with the best accuracy (ADE/FDE of 0.2393/0.3095), AWRS of 0.1528, and a low runtime of 14.94 ms.
TABLE II: Average ADE, FDE, WRS, and runtime(ms) values for Real-Imitated trajectory from three path types [P1, P2, P3] across different algorithms and filter scenarios.
@@
V-D Safe Path Navigation using SafeNavChull and SafeNavCentroid with RiskAware RRT*
We developed two navigation methods called SafeNav-Chull and SafeNav-Centroid by integrating LanBLoc-BMM$(EKF)$ with safety checks based on Algorithms 1 and 2, respectively. Both methods incorporate the proposed RAw-RRT* for obstacle avoidance and risk minimization within the safe path. To evaluate their effectiveness, we conducted trials on three path types formed by sequences of cluster members with different origins and a common destination. The resulting trajectories are shown in Fig. 2 and Fig. 3. Each path type was simulated over $1000$ trials, and performance was evaluated using ADE, FDE, AWRS, percent error, and runtime metrics.
Figure 2: Safe path navigation using SafeNav-Chull and RAw-RRT across three path types $(P1$$P3)$ from different clusters to the common destination (cluster 3). Colored segments represent convex hulls; red and blue lines indicate actual and observed trajectories.
Figure 3: Safe path navigation using SafeNav-Centroid and RAw-RRT across three path types $(P1$$P3)$ from different clusters to a common destination (cluster 3). Red and blue lines represent actual and observed trajectories, respectively.
V-D1 Average Percent Error in Trajectory Lengths
Table III shows the trajectory length error for path types $P1$, $P2$, and $P3$. SafeNav-CHull shows an average error of 6.51%, while SafeNav-Centroid improves this to 5.46%. Notably, P2 has the highest error across all methods, likely due to increased drift caused by noise in velocity and heading factors sensitive to terrain complexity and frequent directional changes.
TABLE III: Path length error of estimated trajectories for each path type

### v2

@@ IV-B3 Position Update:
In this model, the velocity at the next time step, $v_{t+1}$, is updated based on the current velocity, $v_{t}$ constrained by the acceleration/deceleration capabilities and speed limits. The heading update adjusts the heading, $\theta_{t+1}$, by a controlled amount, $\Delta\theta$, within maneuverability limits. The position update calculates the new position, $\vec{p}_{t+1}$, based on the updated velocity and heading. This model assumes a linear motion and instant change of speed and heading for simplicity.
However, in practice, the dynamics of the battlefield are unpredictable due to changing environmental conditions, terrain, and obstacle components. To represent the motion model of the real-battlefield scenario, we need to consider the process noise in position, velocity, acceleration, and heading that capture the uncertainty and variability due to factors like terrain and other unmodeled influences such as obstacles and hazards. It includes terms for terrain effects $(\tau)$ and process noises ${\epsilon}_{x,t},{\epsilon}_{y,t},{\epsilon}_{v_{x},t},{\epsilon}_{v_{y},t},{\epsilon}_{a_{x},t},{\epsilon}_{a_{y},t},{\epsilon}_{\theta,t}$ in as position, velocity, acceleration and heading respectively. Hence, the motion model is defined in Eqns. 3,4, and 5 is updated as:
@@
Terrain effect $\tau$ can significantly influence the error in the predicted position. By accurately computing $\tau$, predictions for the ME’s position can be made more precise, enhancing the navigation strategy in challenging and dynamic terrain conditions. While moving from point P1 $(x1,y1)$ to P2 $(x2,y2)$, total time to travel $t_{\tau}$ considering $\tau$ is given by Eqn 8.
@@
where $\lambda=\tau/v_{c}$ is the time delay due to $\tau$. $\lambda$ is calculated at the beginning with prior knowledge of $\tau$. However, $\tau$ may change as navigating along the field due to changing terrain conditions and the velocity. Thus, the predicted position $p_{t+1}$ of moving entity considering the terrain effect is given by Eqn. 9.
@@
where ${p}_{t}$ is the current position of moving entity.
@@ IV-C Trajectory Estimation using BMM and EKF
The Extended Kalman Filter (EKF) for the non-linear battlefield scenario linearizes the current mean and covariance, allowing it to approximate the state of a nonlinear system, and thus is an effective filtering algorithm due to its low computational complexity and high estimation accuracy. We adapted an EKF by integrating BMM defined in Eqn. 6 and 7 in its prediction step to predict the future state of a moving entity. The mathematical representation of EKF is shown below:
@@
IV-C1 State and Measurement Models
The state of the system at time step $k$ is represented by the vector $x_{k}$. The state transition function $X_{k+1}$ and the measurement function $Z_{k+1}$ are represented in Eqn. 10:
@@ IV-C1 State and Measurement Models
$f(\cdot)$ is the nonlinear state transition function defined in Eqn 9, which predicts the next state based on control input $U_{k}$ and the previous state $X_{k}$ while the $h(\cdot)$ is the LanBloc measurement function, that maps the true state space into the observed space.$W_{k}$ is the process noise and and $V_{k+1}$ is the measurement noise which is normally distributed with covariance $Q_{k}$ and $R_{k+1}$ respectively.
IV-C2 Prediction Step
The EKF predicts the state and error covariance at the time interval $k$. The predicted state estimate $\hat{X}_{k+1}$ and predicted covariance estimate $P_{k+1}$ are given by:
@@ IV-C2 Prediction Step
where $F_{k}$ is the Jacobian of state transition function $f(\cdot)$ with respect to $X$ that is evaluated at $\hat{X}_{k+1}$ and $Q_{k}$ is the covariance matrix of process noise.
IV-C3 Update Step
The EKF incorporates the location measured by the LanBloc algorithm to update the state estimate and covariance. The measurement residual $\tilde{y}_{k+1}$ and residual covariance $S_{k+1}$ are calculated using Eqn. 12.
@@ IV-C3 Update Step
Where $z_{k+1}$ represents the actual measurement of the position at time step $k+1$. The Kalman Gain is a fundamental component of the Kalman Filter that determines the degree to which the new measurement will be incorporated into the state estimate. So, the Kalman Gain $K_{k+1}$ is computed using the predicted covariance estimate $P_{k+1}$, the transpose of the observation model $H_{k+1}^{T}$, and the inverse of the residual covariance $S_{k+1}^{-1}$ in equation 13 below.
@@ IV-D1 Convex Hull Method
The convex hull technique provides a geometrically sound and computationally efficient method to determine the boundaries of a safe path, making it an essential tool in navigation and path-planning applications. Our method utilizes the convex hull approach[13] to confirm the safe zone while making the heading decision toward the look-ahead point while navigating. For safety checks, the path is split into multiple segments and arranged according to their sequence. For each segment, $S=${($x_{1}$,$y_{1}$),($x_{2}$,$y_{2}$),…,($x_{n}$,$y_{n}$)}, convex hull is calculated using Chan’s algorithm [14] and the predicted state is checked against the convex hull boundary exploiting the point in polygon test [15].
At the beginning of each trial, the node obtains its initial position $(x_{i},y_{i}$ using LanBLoc while velocity $(v_{i})$ and heading $(\theta_{i})$ is measured using IMU sensor simulator which serves as an input to EKF module that predicts the next state $(x_{P},y_{P},v_{P},{\theta}_{P})$ based on the motion model discussed in detail in Section IV-B. Then, the node checks if the predicted position lies within the convex hull indicating the safe zone, and makes the heading decisions. If the position lies within the convex hull of the current or next segment, it moves with the default command (desired speed, desired heading). Otherwise, it re-calculates the control commands. Node steps back with the default control command if the predicted position is not within any of the adjacent hulls(condition when a node is in the boundary of the safe path.) The trajectory prediction mechanism is explained in detail in Section IV-C while the safety check using a convex hull is discussed in Algorithm 1.
@@ IV-D2 Centroid Method
The Centroid Method relies on the centroids of sub-segments within a safe path to compute the control commands. In this approach, the current path segment is divided into smaller sub-segments, and the centroids of these sub-segments are calculated. The algorithm determines the closest centroid to the entity’s current position(updated by EKF) and computes control input $U(v_{i},{\theta_{i}})$ required to reach the centroid. Using these control commands, it predicts the next position of the entity which is mentioned in detail in the Algorithm 2. This algorithm attracts the moving entity toward the centroid of the path segment with an objective of minimizing the drift from the central trajectory.
@@ IV-E1 Rapidly Exploring Random Tree-Star (RRT*)
RRT* [16] relies only on the path length to select the optimal trajectory. During the optimization phase RRT* searches to find the shortest path between consecutive nodes $\mathbf{x}_{i}$ and $\mathbf{x}_{i+1}$ using the cost function in Eqn 15.
@@
IV-E2 Risk-Aware RRT*
We propose a new variant of RRT* called Risk-Aware RRT* (RAw-RRT*) modifying its cost function in Eqn. 15 which ensures the obstacle-free optimal path with the shortest length and minimal risk of hazards. We use this algorithm for re-generating the path that avoids obstacles along the predicted next position while running CHull and Centroid-based Controlled maneuvering. For a trajectory with points $\{\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{n}\}$, the Risk-Aware cost function in RAw-RRT* is defined as:
@@ IV-E2 Risk-Aware RRT*
Where, $C_{\text{total}}$ is the Total cost of a trajectory, $C_{\text{length}}$ is the total path length calculated as the sum of Euclidean distances between consecutive points. $C_{\text{risk}}$ is the total risk cost, calculated as the sum of Weighted Risk Scores (WRS) for each point $\mathbf{p}_{i}$ in the trajectory. $\alpha,\beta$ are the weight parameters that control the trade-off between path length and risk.
@@ IV-E3 Edge-Wise WRS Calculation
For each edge in the trajectory, the Weighted Risk Score (WRS) is calculated based on the deviation of the edge from the central trajectory. Let $\mathbf{c}_{i}$ represent the central trajectory point closest to $\mathbf{p}_{i}$, then:
A. Deviation:
@@ B. Zone-Based Weighted Risk Score :
Where, $|d_{i}|$ is the absolute deviation for point $\mathbf{p}_{i}$. $a_{j}$ and $b_{j}$ are the lower and upper bounds of zone $Z_{j}$ and $w_{j}$ is the maximum risk weight of $Z_{j}$ $d_{\text{max}}$ is the maximum deviation and $w_{\text{max}}$ is the maximum weight assigned when $|d_{i}|\geq d_{\text{max}}$. For $|d_{i}|<d_{\text{max}}$, the Eqn. 18 computes the weighted risk score based on the absolute deviation and the zone-specific weight $w_{j}$.
For $|d_{i}|\geq d_{\text{max}}$, the deviation is capped at $d_{\text{max}}$ and a maximum weight $w_{\text{max}}$ is used to calculate the risk score.
This ensures that both small and large deviations are treated appropriately, with large deviations capped at $d_{\text{max}}$ and assigned a constant maximum risk score.
IV-E4 Total Cost for Trajectory
For a trajectory with points $\{\mathbf{p}_{i}..\mathbf{p_{n}}\}$, the total risk-aware cost is given by Eqn. 19.
$C_{\text{total}}=\alpha\cdot\sum_{i=1}^{n-1}\|\mathbf{p}_{i}-\mathbf{p}_{i+1}\|+\beta\cdot\sum_{i=1}^{n}\text{WRS}(\mathbf{p}_{i})$ (19)
The detailed steps to obtain optimal path using risk-aware cost function are systematically presented in Algorithm 3.
Algorithm 3 Risk-Aware RRT*
1:Start position $p_{\text{start}}$, goal position $p_{\text{goal}}$, obstacle map $O$, risk weights $w_{i}$
2:Optimal path $P_{\text{optimal}}$
3:Initialize tree $T\leftarrow\{p_{\text{start}}\}$, set cost $C(p_{\text{start}})\leftarrow 0$
4:for $k=1$ to $N_{\text{max}}$ do
5:     Sample $p_{\text{rand}}$, find $p_{\text{nearest}}\in T$, and steer to $p_{\text{new}}$
6:     if $\text{CollisionFree}(p_{\text{nearest}},p_{\text{new}},O)$ then
7:         Compute: $C_{\text{total}}(p_{\text{new}})=C_{\text{total}}(p_{\text{nearest}})+\|p_{\text{nearest}}-p_{\text{new}}\|+\text{WRS}(p_{\text{new}})$
8:         Add $p_{\text{new}}$ to $T$ and rewire nearby nodes for cost optimization
9:     end if
10:end for
11:Extract optimal path $P_{\text{optimal}}$ from $T$
12:return $P_{\text{optimal}}$
@@ IV-F Safe Path Navigation using SafeNav-Chull and SafeNav-Centroid and RAw-RRT*
9:         Run Algorithm 3 to find alternative path to nextPosition that minimize cost.
@@ V Experiments and Results
To test the validity of our proposed navigation algorithms, experiments were performed on Alienware Aurora R12 System with 11th Gen Intel Core i7 CPU, 32 GiB Memory, and Nvidia GeForce RTX 3070 GPU using Python 3.11 on PyCharm CE 2024 IDE. The landmark recognition model was trained using Python-3.10.12 and torch-2.5.1+cu121, utilizing NVIDIA A100-SXM4 GPU with 40GB memory.
@@ V-A Datasets
V-A1 MSTlandmarkv1 Dataset
We have used publicly available MSTlandmarkv1[3] dataset to train the yolov11-based landmark recognition model. The dataset was split into $70:20:10$ for training, validation, and testing. Image augmentation (brightness adjustment, bounding box rotation & noise addition, was done on the training portion of the dataset to avoid overfitting problems and to improve the model’s robustness.
V-A2 MSTlandmarkStereov1 Dataset
We used a real-world landmark stereo dataset[3] for the node-to-landmark distance estimation. We used two identical cameras of $(2560*1440px~{}@~{}30~{}fps)$ resolution with an adjustable field of view (FoV) to collect this dataset. The cameras were arranged on the metal bar with adjustable baseline(B) from $(10-40)$ centimeters. This dataset was used to perform distance estimation from unknown nodes to the landmark anchors.
V-A3 Trajectory Dataset and Safe Path
@@
ii) Real Trajectory Dataset: The real trajectory and safe path datasets were obtained to validate the proposed navigation framework. We set a battlefield region of $200*200$ grid space and mapped all the real-world geographical landmarks from MSTlandmarkStereov1 datasets. We randomly assigned 25 obstacles and 25 hazard components within the grid space to simulate the risk and obstacle scenario in the real battlefield as shown in Fig. 1. Using the landmark dataset, clusters with at least three landmarks within detectable range are selected as the trilateration set for LanBloc localization. Path clusters, defined as sequences of landmark clusters, are used to generate ground truth trajectories from start to end clusters, based on movement through cluster centroids. These trajectories are represented as polylines connecting cluster centroids. The object’s states along these trajectories are considered the ground truth trajectory, with trajectory points represented by seven state variables, similar to the synthetic trajectory.
iii) Safe Path Buffer and Risk Zone: A safe buffer is a bounded region defined on both sides of the ground truth trajectory as in Fig.1. This bounded region ensures safe navigation and is divided into four equal-width zones. we considered $w_{1}=2,w_{2}=4,w_{3}=6,w_{4}=8$ as the maximum risk weights assigned to each zone on either side of the central trajectory. Weighted Risk score anywhere within a zone is calculated based on deviation from the ground truth trajectory using Eqn 18 which increases with distance from the central trajectory and quantifies risk exposure.
@@ V-B Evaluation Metrics
Standard evaluation metrics such as Average Displacement Error (ADE), Final Displacement Error (FDE), and Percent Error were used for estimated trajectory evaluation. We also introduced a new metric called Average Weighted Risk Score(AWRS) to evaluate an overall weighted risk exposure of estimated trajectories. To ensure robust and reliable results, the object’s motion within each safe path was simulated for 1000 trials, and their average values were calculated. The Average Weighted Risk Score (AWRS) is the mean of the individual Weighted Risk Scores across all trajectory points shown in Eqn. 20.
$\text{AWRS}=\frac{1}{N}\sum_{i=1}^{N}\text{WRS}_{i}$ (20)
Where N is the total number of trajectory points and $\text{WRS}_{i}$ is the weighted risk score for the $i^{th}$ point calculated using Eqn. 18.
$ADE=\frac{1}{N}\sum_{t=t_{0}}^{t_{f}}{\|{{\hat{Y}_{(t)}-{Y}_{(t)}}}\|}$ (21)
The ADE is an average of pointwise L2 distances between the estimated and ground truth trajectory calculated using Eqn. 21.
The FDE is the L2 distance between the final points of the estimated and ground truth trajectory using Eqn 22.
$FDE={\|{{\hat{Y}_{\left({t_{f}}\right)}-{Y}_{\left({t_{f}}\right)}}}\|}$ (22)
where $\hat{Y}_{t}$ is the predicted position at time step t and $Y_{t}$ is the ground truth position. The % error of the estimated trajectory indicates a relative error in path length between the predicted trajectory and the actual trajectory calculated using Eqn. 23.
Percent Error=|Est. LengthTrue Length|True $\text{Percent Error}=\frac{|\text{Est. Length}-\text{True Length}|}{\text{TrueLength}}\times 100\%$ (23)
@@ V-C Simulation and Experimental Results Discussion
Baselines: To evaluate the effectiveness of the proposed navigation algorithm (LanBLoc-BMM), we compared it with various state-of-the-art visual localization and Multi-Source Fusion methods from literature such as DeepLoc-GPS, DeepLoc-EKF-GPS [8], Multiple Sensor Fusion [19] using Synthetic and Real-datasets.
The performance results of four algorithms combined with EKF and PF using Synthetic datasets for ground navigation are shown in Table II and using real datasets in Table III. Table II highlights the best overall performance of LanBLoc(EKF), with the lowest ADE, FDE, AWRS, and runtime, while DeepLoc-GPS-EKF (EKF) is a close contender, performing well in all metrics, particularly in FDE and AWRS. DeepLoc-GPS (EKF) also performs well, especially in ADE and FDE; however, it lags in runtime. Sensor Fusion (EKF), while having the highest ADE, shows competitive FDE but is less efficient in computation.
TABLE II: Performance across Algorithms on synthetic trajectory data
@@
From Table III, it is evident that LanBLoc-BMM (EKF) is best regarding accuracy, weighted risk score, and runtime for real trajectory datasets. It consistently offers the most accurate results with reasonable computation times and minimal risk exposure, making it the preferred choice for accurate and efficient localization across different path types. DeepLocGPS(EKF) provides its best accuracy with an ADE/FDE of 0.3041/0.3103, AWRS of 0.2009, and a reasonable runtime of 20.58ms. DeepLocGPSEKF shows the best accuracy with an ADE/FDE of 0.2869/0.2969, AWRS of 0.1906, and a runtime of 18.94 ms. Sensor Fusion(EKF) provides a balance of accuracy (ADE/FDE of 0.2998/0.3115), AWRS of 0.1749, and an acceptable runtime of 22.31ms. The proposed LanBLoc-BMM (EKF) based navigation again stands out with the best accuracy (ADE/FDE of 0.2393/0.3095), AWRS of 0.1528, and a low runtime of 14.94ms.
TABLE III: Average ADE, FDE, WRS, and runtime(ms) values for Real trajectory from three path types [P1, P2, P3] across different algorithms and filter scenarios.
@@
Figure 3: Comparison of Weighted Risk Exposure of EKF-based Algorithms. Fig 3.1(a) and 3.1(b) show the overall WRS performance of four localization algorithms while using EKF and PF, respectively. The Fig 3.2(a),3.2(b),3.2(c) and 3.2(d) shows the individual comparison of weighted risk score(WRS) for EFK and PF integration across each of the four localization algorithms.
Fig. 3 shows the overall and individual comparison of WRS of trajectories generated by four different algorithms while integrating EKF and PF separately. Fig. 3.1(a) compares the WRS of trajectories generated by four different algorithms when EKF is applied while Fig. 3.1(b) shows a similar comparison while PF is applied. Also, Fig. 3.1(a),(b),(c), and (d) shows the WRS comparison of each algorithm while applying EKF and PF separately.
V-D Safe Path Navigation using SafeNav-Chull and SafeNav-centroid with Risk-Aware RRT*
Next, we design two navigation methods integrating LanBLoc-BMM(EKF) called SafeNav-CHull and SafeNav-Centroid, combining two safety check approaches in Algorithms 1 and 2 respectively for safety checks and controlled maneuvering. Then, we integrated the novel RAw-RRT* algorithm with these two navigation approaches for obstacle avoidance and risk minimization within the safe path. To test the efficacy of these two navigation algorithms, We ran the navigation trials on three path types drawn from the combination of consecutive cluster members starting from different origins and reaching the same destination. Then, we obtained the trace of the positions that an object traveled along to reach the destination as shown in Fig. 4 and Fig. 5. We simulated 1000 thousand movement trials of an object along each path type and averaged their performance in ADE/FDE, AWRS, and percent error and runtime.
V-D1 Average Percent error in trajectory lengths of two approaches
Table IV shows the % error in the length of the estimated and ground truth trajectory for different path types: $P1$, $P2$, and $P3$. We obtained an average error of $6.51\%$ for SafeNav-CHull. For the same three path types, we obtained an improved average error of $5.46\%$ for SafeNav-Centroid.
Figure 4: Showing Safe Path Navigation using SafeNav-Chull and RAw-RRT* in three different path types obtained from the combination of different landmark clusters and starting points that guide to the common destination (cluster 3). Fig. 4 (a) shows the path type 1 (P1) from the centroid of cluster 6 to cluster 3. Fig. 4 (b) shows the path type 2 from cluster 4 to cluster 3, and Fig. 4 (c) shows the path type 3 (P3) from cluster 7 to cluster 3. The color coding on each path type represents the convex hull of the path segments. The actual and observed trajectories of moving entities within each path type are shown by red and blue lines, respectively.
Figure 5: Showing Safe Path Navigation using SafeNav-Centroid and RAw-RRT* in three different path types obtained from the combination of different landmark clusters and starting points that guide to the common destination (cluster 3). Fig. 5 (a) shows the path type 1(P1) from the centroid of cluster 6 to cluster 3. Fig. 5 (b) shows the path type 2 from cluster 4 to cluster 3 and Fig. 5 (c) shows the path type 3 (P3) from cluster 7 to cluster 3. The actual and observed trajectories of moving entities within each path type are shown by red and blue lines, respectively.
From the % error observed across the three path types, it is evident that the trajectory of the moving object in P2 exhibits higher errors for all algorithms compared to P1 and P3. This increased error is likely due to the drift in the estimated trajectory, which is positively correlated with noise in state variables such as velocity and heading direction. These variables are particularly sensitive to the terrain’s characteristics, shape, and length of the path. When an object traverses a path with many turns and challenging terrain, velocity and heading need to be measured more frequently, potentially leading to greater error accumulation than on a path with fewer turns and bends.
TABLE IV: Path length error of estimated trajectories for each path type